OpenStudy (anonymous):

Sasha buys two ribbons of equal length. She cuts 9 equal pieces from the first ribbon and 5 inches are left over. She cuts two pieces of the same length from the second ribbon and 40 inches are left over. What is the length of each ribbon?

hero (hero):

x = 1st ribbon y = 2nd ribbon x - 9b = 5 y - 2b = 40 x = 9b + 5 y = 2b + 40 x = y 9b + 5 = 2b + 40 9b - 2b = 40 - 5 7b = 35 b = 35/7 b = 5 x = 9(5) + 5 x = 45 + 5 x = 50 y = 2(5) + 40 y = 10 + 40 y = 50 Therefore each ribbon is 50 inches long.